# Unveiling the Potential of iMarkers: Invisible Fiducial Markers for Advanced Robotics

### v1

@@ 5.1 Operational Implementation
Software. Based on the detection algorithms outlined in Section 4, two software variants have been developed to support iMarker detection for general and robotics researchers. These variants include a standalone Python application with a Graphical User Interface (GUI) for visualization and configuration and a Robot Operating System (ROS)-based version designed for robotics tasks. Both variants will be publicly available soon at https://snt-arg.github.io/iMarkers/.
@@ 5.3 Detection Range Evaluation
A comparative experiment was conducted using printed markers and iMarkers, both formatted as $7\times 7~{}\mathrm{cm}$ ArUco markers [11], to evaluate the effectiveness of iMarkers w.r.t. traditional markers and analyze potential variations in detection performance. The markers were affixed to a vertical surface, and detection tests were performed from various distances and viewing angles. The objective was to identify the maximum distance at different angles where the markers remained detectable using the ArUco detection algorithm. The experiments utilized a static single-vision setup (shown in Fig. 6) for iMarker detection, while the same sensor without a polarizer was used for printed marker detection.
Figure 8: Maximum distance required to detect “printed” and “iMarkers” ArUco formats in various angles of view.
According to the experimental results depicted in Fig. 8, the static and dynamic single-vision setups show negligible differences in maximum detection ranges, differing only by an average of $5~{}cm$. However, the disparity in detection ranges between $7\times 7~{}\mathrm{cm}$ printed markers and iMarkers is more pronounced, ranging from $15$ to $40~{}cm$. This variation is minor when the viewing angle approaches $75^{\circ}$ and more significant for more direct angles ($0^{\circ}$). Considering the purpose of introducing iMarkers, and given that printed markers are specifically designed as high-contrast landmarks for optimal visibility, such results are expected. iMarkers offer a distinct advantage with their unobtrusive design, allowing unnoticeable integration into environments without compromising visual aesthetics. This trade-off highlights the practicality of iMarkers for applications where invisibility to the human eye is crucial while maintaining reliable detection by robotic systems.
@@ 5.4 Recognition Speed
GUI-related $71~{}\mathrm{ms}$ $67~{}\mathrm{ms}$ $69~{}\mathrm{ms}$ $71~{}\mathrm{ms}$ $72~{}\mathrm{ms}$
@@
Post-processing $3~{}\mathrm{ms}$ $4~{}\mathrm{ms}$ $5~{}\mathrm{ms}$ $4~{}\mathrm{ms}$ $5~{}\mathrm{ms}$
Recognition $2~{}\mathrm{ms}$ $2~{}\mathrm{ms}$ $3~{}\mathrm{ms}$ $2~{}\mathrm{ms}$ $4~{}\mathrm{ms}$
Sum $92~{}\mathrm{ms}$ $118~{}\mathrm{ms}$ $128~{}\mathrm{ms}$ $162~{}\mathrm{ms}$ $178~{}\mathrm{ms}$
@@
Considering the necessity of fast recognition speeds in demonstrating the practicality of fiducial markers for real-time applications, this subsection provides a detailed profiling analysis of iMarker detection and recognition. Table 2 summarizes cumulative processing times for iMarker data decoding using various sensor configurations discussed in subsection 5.1. Evaluations were conducted using the implemented standalone Python application introduced in the mentioned subsection. To obtain the measurements, $7\times 7~{}\mathrm{cm}$ iMarkers were placed approximately $50~{}\mathrm{cm}$ from the lens, directly facing the sensor under normal illumination conditions. For the UV-range iMarker, a near-UV light source was also employed.
According to the table, the overall iMarker recognition process, from image acquisition and user interface loading to marker recognition, is completed within milliseconds. To compare, some methods discussed in Table 1 require several seconds for detection, highlighting the efficiency of the iMarker system. In dual-vision setups, frame acquisition and processing account for the subsequent significant portions of the processing time due to the need to receive synchronized frames from two cameras and perform image alignment for further processing, respectively. However, the post processing and marker recognition procedures are executed exceptionally quickly, owing to the binary nature of the image feeds.
@@ 6.1 Comparative Insights
Fig. 9 depicts relative performance scores of different iMarker fabrication variants and the designed sensor configurations for their detection. Accordingly, UV-range markers exhibit the highest level of invisibility, detection equipment cost (sensor, flashlight, etc.), and sensitivity to ambient light. This variant of iMarkers is ideal for applications demanding a high degree of constraint and security. In addition, the user should orient the near-UV light source toward the iMarker, which might be challenging due to the marker’s invisible nature. Visible-range camouflaged iMarkers achieve ideal readability while remaining minimally visible to the human eye upon close inspection. The advantages of this variant are a more straightforward sensor design and increased detectability without additional illumination. Finally, despite maintaining appropriate unobtrusiveness, the IR-range version shows more vulnerability to environmental factors due to potential occlusion by materials that absorb or block infrared light. IR-range iMarkers also demand a an IR-sensitive sensor and illuminator for detection.
@@
Regarding the sensors, dual-vision setups are vulnerable to mechanical component misalignment (i.e., camera and beamsplitter), resulting in noisy frame subtraction and reduced marker detection accuracy. Unlike the dual-vision cube beamsplitter setup with standard optical cage cubes, the plate beamsplitter variant is more susceptible to misalignment due to its custom-designed mount. Hence, vibration and thermal expansion can cause the screws to become loosened or the cameras to shift out of their precise alignment, reducing the reliability of the captured data. Dual-vision setups also restrict the cameras’ FoV by excluding regions outside the beamsplitter coverage, regardless of its size or variation. However, they demand the least computational cost for iMarker detection by filtering out non-CSR regions and processing only the output binary image. Static single-vision setup is the most uncomplicated sensor design, imposing higher algorithmic overhead for iMarker detection. This solution may encounter challenges in scenarios with varying illumination, necessitating adjustments in “masking” and “threshold” parameters. Moreover, single-vision setups are more feasible for minimizing payloads, such as in drones, as dual-vision configurations are heavier. The combined weight of dual-vision setup components typically ranges from 150 to 350 grams, with the plate beamsplitter potentially reducing the overall sensor weight.
@@ 6.2 Applicability
Considering the unique characteristics of iMarkers, they serve as potential alternatives in scenarios reliant on printed markers. By safely replacing printed markers, iMarkers can preserve unobtrusiveness and aesthetic integrity while delivering full functionality for robotics and AR systems. Their obscure nature makes them appropriate for surroundings with unwanted visual clutter, such as museums and offices. Fig. 10 depicts the real-world benchmarking of iMarkers in visual Simultaneous Localization and Mapping (SLAM), showcasing the use of a visible-range iMarker to facilitate indoor map reconstruction enriched with semantic information [20]. iMarkers can also be suitable for autonomous robots functioning in retail stores and smart warehousing, where numerous markers are needed to augment objects. Another potential use case includes object grasping in cluttered environments, where iMarkers can supply accurate localization cues to enhance robotic manipulation accuracy. iMarkers can also be strategically employed to seamlessly augment difficult-to-detect objects, like transparent surfaces (e.g., windows and glass block walls), where the transparency poses considerable challenges to computer vision and robot perception systems. Likewise, reflective surfaces, such as mirrors or chrome/metallic objects, often cause distortions or reflections, confusing perception algorithms and negatively impacting object recognition or localization accuracy. In these scenarios, iMarkers can augment such objects by providing reliable information while preserving visual appeal. This allows robots and AR devices to effectively see and interact with visually challenging entities, ensuring reliable operation and versatility.
@@ 6.3 Limitations
Considering the ongoing nature of the research project and the technical fabricating process of CSR shells, the experiments are constrained by the limited variety and quantity of available shells, restricting the ability to produce and evaluate many iMarkers. Additionally, long-term exposure to environmental factors may affect marker visibility, as markers tend to develop a yellowish tint over time.
@@ 7 Conclusions
This paper introduced a novel generation of fiducial markers, termed iMarkers, designed explicitly for real-world robotics applications. iMarkers are designed to be visually indistinguishable from the human eye, ensuring minimal environmental disruption while remaining fully detectable and recognizable by robots equipped with specialized sensors and advanced detection algorithms. The paper also presents various sensor designs and implemented algorithms for detecting iMarkers placed in the environment and exploiting them for different use cases. Experimental results demonstrate the practicality and potential of these markers as versatile landmarks for data embedding in various real-world scenarios.
Future works include optimizing the iMarker fabrication process to ensure scalable production and enhancing sensor designs and algorithm implementations to guarantee robustness across various environmental conditions. Another plan is to conduct visual SLAM experimentation in real-world environments augmented with iMarkers.

### v2

@@ 5.1 Operational Implementation
Software. To support the detection algorithms outlined in Section 4, an open-source iMarker detection software is developed, publicly available at https://snt-arg.github.io/iMarkers/. The framework is implemented in Python and features a Robot Operating System (ROS) interface, enabling straightforward integration into robotic applications.
@@ 5.3 Detection Range Evaluation
A comparative experiment was conducted using printed markers and iMarkers, formatted as ArUco library [11], to evaluate the effectiveness of iMarkers w.r.t. their classic counterparts. The markers were affixed to a tripod with an adjustable mounting plate, enabling precise control over the viewing angle relative to the camera. The objective was to measure the maximum distance across different viewing angles where each marker variant remained detectable using the ArUco detection algorithm. Accordingly, an Intel® RealSense™ D435 camera was used across all experiments: unmodified for printed markers and equipped with static or dynamic polarizers for iMarker detection, ensuring consistent sensor characteristics for fair comparison. The examined viewing angle ranged from $0^{\circ}$ to $75^{\circ}$, within which the estimated poses of the markers remained reliable. Experiments were conducted under normal lighting conditions, and $7\times 7~{}\mathrm{cm^{2}}$ visible-range iMarker variants were used to be compared against $6\times 6~{}\mathrm{cm^{2}}$ and $7\times 7~{}\mathrm{cm^{2}}$ printed ArUco markers as references. Experimental results are depicted in Fig. 8.
Figure 8: Maximum detection distances for “printed” and “iMarkers” ArUco formats at various viewing angles. The shaded regions (error bands) indicate measurement uncertainty in maximum recognition distance.
According to the evaluation results, increasing the viewing angle reduces the maximum marker detection range in all scenarios, mainly due to perspective distortion. Additionally, iMarker detection using static and dynamic single-vision setups shows negligible differences in maximum detection ranges, differing between $3~{}\mathrm{cm}$ to $7~{}\mathrm{cm}$, with an average of $4.8~{}\mathrm{cm}$. However, regardless of the detection setup, printed markers achieved consistently higher detection ranges than iMarkers of identical dimensions ($7\times 7~{}\mathrm{cm^{2}}$). The detection range of $7\times 7~{}\mathrm{cm^{2}}$ iMarkers closely matches that of $6\times 6~{}\mathrm{cm^{2}}$ printed ArUco markers, with differences ranging from $4~{}\mathrm{cm}$ to $6~{}\mathrm{cm}$ and averaging $4.2~{}\mathrm{cm}$ across various angles of view. It should be noted that printed ArUco markers, thanks to their high-contrast black-and-white patterns, offer substantial visual features that facilitate detection. However, iMarkers demonstrate comparable detection performance in real-world experiments, apart from their distinct advantage over classic markers: their ability to seamlessly blend into the environment without disrupting visual aesthetics. This makes iMarkers highly suitable for shorter-range applications where visual discretion is essential, such as indoor navigation.
@@ 5.4 Recognition Speed
Post-processing $3~{}\mathrm{ms}$ $9~{}\mathrm{ms}$ $8~{}\mathrm{ms}$ $4~{}\mathrm{ms}$ $5~{}\mathrm{ms}$
Recognition $2~{}\mathrm{ms}$ $2~{}\mathrm{ms}$ $3~{}\mathrm{ms}$ $2~{}\mathrm{ms}$ $3~{}\mathrm{ms}$
Sum $21~{}\mathrm{ms}$ $56~{}\mathrm{ms}$ $62~{}\mathrm{ms}$ $91~{}\mathrm{ms}$ $105~{}\mathrm{ms}$
@@
Considering the necessity of fast recognition speeds in demonstrating the practicality of fiducial markers for real-time applications, this subsection provides a detailed profiling analysis of iMarker detection and recognition. Table 2 summarizes cumulative processing times for iMarker data decoding. Evaluations were conducted using various sensor configurations and implemented software introduced in subsection 5.1. The algorithms developed for all sensor variants share five core modules: acquisition,” which captures video frames; preprocessing,” which applies image enhancement techniques; processing,” which refines frames to emphasize marker patterns; post-processing,” which enhances the final binary images; and recognition to detect and interpret marker patterns. To obtain the measurements, $7\times 7~{}\mathrm{cm^{2}}$ iMarkers were placed approximately $50~{}\mathrm{cm}$ from the lens, directly facing the sensor under standard lighting conditions with evenly diffused artificial illumination. For the UV-range iMarker, a DARKBEAM A300 $365~{}nm$ flashlight was additionally used to enhance the iMarker visibility.
According to Table 2, the overall iMarker recognition process, from video frame acquisition to inner pattern recognition steps, is completed within milliseconds. To compare, some methods discussed in Table 1 require several seconds for detection, highlighting the efficiency of the iMarker system. Among the various configurations, UV-range iMarker recognition (Algorithm 4) is the fastest, as the input image is already in grayscale, requiring only a simple “thresholding” operation to reveal the marker pattern for recognition. In dual-vision setups (Algorithm 1), acquisition and processing account for the subsequent significant portions of the processing time due to the need to receive synchronized frames from two cameras and perform image alignment for further processing, respectively. The post-processing stage is executed more efficiently, as the substantial filtering process retains only the differing image regions, i.e., the CSR areas. On the other hand, single-vision setups operate faster in acquisition due to capturing frames from a single camera. However, they allocate more time to post-processing, as the captured images contain higher noise levels and the need to refine iMarker regions effectively.
@@ 6.1 Comparative Insights
Fig. 9 depicts relative performance scores of different iMarker fabrication variants and the designed sensor configurations for their detection, where larger areas indicate higher values. The relative performance scores were assigned based on theoretical analysis and experimental observations, which will be described in detail.
In Fig. 9-a, three primary performance aspects of the iMarker variants have been analyzed. Regarding “invisibility,” UV-range iMarkers remain entirely imperceptible to the eye, while IR-range iMarkers may exhibit slight visibility due to scattering under extreme lighting conditions. Visible-range iMarkers, on the other hand, are designed to blend into patterned backgrounds; however, upon close inspection, they remain minimally visible to the human eye. When “readability” is studied, visible-range iMarkers have better compatibility with standard cameras and do not require additional illumination. However, IR- and UV-range iMarkers require dedicated illumination sources and particular sensors, with varied detection performance based on environmental conditions. Finally, and in terms of “detection equipment cost” (camera, flashlight, etc.), visible-range iMarkers are the most economical, as they can be detected using standard cameras. UV equipment (camera and illumination sources) is generally the most expensive among the three, and IR-range iMarker detection equipment cost lies in between.
@@
Considering the sensor-related factors shown in Fig. 9-b, the performance scores can be studied in five aspects. In “design complexity,” the static single-vision setup is the most straightforward, requiring only a static polarizer attached to the lens. The dynamic single-vision setup has a moderate complexity by incorporating a function generator to control polarization. Dual-vision configurations represent the most complex designs, with the cube beamsplitter offering more reliable light division than the plate one, which requires precise alignment for optimal beam splitting. The design complexity directly influences the “sensor cost.” The static single-vision setup offers the most economical solution, requiring only a polarizer, while the dynamic single-vision configuration incurs additional costs due to its polarization control components. Dual-vision setups are generally more expensive. They require two cameras, a beamsplitter (either cube or plate), and a structural mounting cage to ensure precise alignment and stability.
Regarding the “field of view,” single-vision setups provide a broader theoretical FoV (e.g., $87^{\circ}\times 58^{\circ}$ in the RealSense™ D435, shown in Fig. 6-c). However, in dual-vision setups, the beamsplitter’s physical dimensions and the supporting cage partially obstruct the cameras’ FoV, reducing the effective viewing area. For instance, the configuration shown in Fig. 6-c achieves an effective FoV of $57^{\circ}\times 54.5^{\circ}$. Another essential factor is “weight”; the static single-vision setup is the lightest, whereas the dynamic single-vision setup adds approximately $30~{}gr$ due to the function generator and dynamic polarizer. Thus, single-vision setups are more feasible for minimizing payloads, such as in drones, as dual-vision configurations are heavier. The combined weight of dual-vision setup components typically ranges from $150$ to $350~{}gr$, with the plate beamsplitter potentially reducing the overall sensor weight. Finally, considering the noise levels in the processed output, dual-vision setups produce clearer results, isolating only the CSR regions of the iMarkers. In contrast, the dynamic single-vision setup may introduce additional noise due to motion and illumination fluctuations during “temporal” subtraction. The static single-vision setup retains the highest noise pixels (i.e., pixels that do not contribute to iMarker detection), including the filtered areas with the detected iMarkers.
@@ 6.2 Applicability
Considering the unique characteristics of iMarkers, they serve as potential alternatives in scenarios reliant on printed markers. By safely replacing printed markers, iMarkers can preserve unobtrusiveness and aesthetic integrity while delivering full functionality for robotics and AR systems. Their obscure nature makes them appropriate for surroundings with unwanted visual clutter, such as museums, hospitals, offices, etc. Fig. 10 depicts the real-world benchmarking of iMarkers in Visual Simultaneous Localization and Mapping (VSLAM), showcasing the use of a visible-range iMarker to facilitate indoor map reconstruction enriched with semantic information [20]. iMarkers can also be suitable for autonomous robots functioning in retail stores and smart warehousing, where numerous markers are needed to augment objects. Another potential use case includes object grasping in cluttered environments, where iMarkers can supply accurate localization cues to enhance robotic manipulation accuracy. iMarkers can also be strategically employed to seamlessly augment difficult-to-detect objects, like transparent surfaces (e.g., windows and glass block walls), where the transparency poses considerable challenges to computer vision and robot perception systems. Likewise, reflective surfaces, such as mirrors or chrome/metallic objects, often cause distortions or reflections, confusing perception algorithms and negatively impacting object recognition or localization accuracy. In these scenarios, iMarkers can augment such objects by providing reliable information while preserving visual appeal. This allows robots and AR devices to effectively see and interact with visually challenging entities, ensuring reliable operation and versatility.
@@ 6.3 Limitations
Given the ongoing nature of this research and the absence of a rapid iMarker fabrication process so far, the experiments remain constrained under varying illumination, weather conditions (for outdoor applications), and environmental factors (e.g., dirt). Additionally, the reflectivity of UV-range CSRs needs to be enhanced to ensure that all UV-range iMarkers within the FoV are reliably detected when illuminated by a broad-field light source.
@@ 7 Conclusions
This paper introduced a novel generation of fiducial markers, termed iMarkers, designed explicitly for real-world robotics applications. iMarkers are designed to be visually indistinguishable by the human eye, ensuring minimal environmental disruption while remaining fully detectable and recognizable by robots equipped with adequate sensors and adapted detection algorithms. The paper also presents various sensor designs and implemented algorithms for detecting iMarkers placed in the environment and exploiting them for different use cases. Experimental results demonstrate the practicality and potential of these markers as versatile landmarks for data embedding in various real-world scenarios.
Future works include optimizing the iMarker fabrication process to ensure scalable production and enhancing the illumination equipment, sensor designs, and algorithm implementations to guarantee robustness across various scenarios.